OpenStudy (anonymous):

2y − x + 2 = 0 -3x + 8y = -5 (-10, 4) (3, 1/2) (2, 0) (4, 1)

OpenStudy (anonymous):

MULTIPLY EQUATION 1 BY 3 ND EQUATION 2 BY 1... SO YOU WILL GET: -3x+6y+6=0 -3x+8y+5=0 subtracting these you will get -2y+1=0 y=1/2 now x can be find out by putting y in equation 1..:)